Sodium hydrogen carbonate (NaHCo), also known as sodium bicarbonate or baking soda, can be used to relieve acid indigestion. Acid indigestion is the burning sensation you get in your stomach when it contains too much hydrochloric acid (HCI) , which the stomach secretes to help digest food. Drinking a glass of water containing dissolved NaHCO3 neutralizes excess HCl through this reaction: HCl(a) NaHCO3(aq) - NaCl(aq)H20) CO29) alo The CO2 gas produced is what makes you burp after drinking the solution Suppose the fluid in the stomach of a woman suffering from indigestion can be considered to be 100. mL of a 0.064 M HCl solution. What mass would she need to ingest to neutralize this much HCl? Be sure your answer has the correct number of significant digits. Periodic Table
Answer #1

lets calculate the mol of HCl

volume , V = 1*10^2 mL

= 0.1 L

use:

number of mol,

n = Molarity * Volume

= 6.4*10^-2*0.1

= 6.4*10^-3 mol

According to balanced equation

mol of NaHCO3 reacted = (1/1)* moles of HCl

= (1/1)*6.4*10^-3

= 6.4*10^-3 mol

This is number of moles of NaHCO3

Molar mass of NaHCO3,

MM = 1*MM(Na) + 1*MM(H) + 1*MM(C) + 3*MM(O)

= 1*22.99 + 1*1.008 + 1*12.01 + 3*16.0

= 84.008 g/mol

use:

mass of NaHCO3,

m = number of mol * molar mass

= 6.4*10^-3 mol * 84.01 g/mol

= 0.5377 g

Answer: 0.538 g